If a system has 425 kcal of work done to it, and releases 5.00×102 kJ of heat into its surroundings, what is the change in internal energy (Δ? or Δ?) of the system?
Answer #1

W = 425 kcal
= 425*4.184 KJ
= 1778.2 KJ

Q = -500 KJ

Use;
Δ? = Q + W
= -500 KJ + 1778.2 KJ
= 1.28*10^3 KJ

Answer: 1.28*10^3 KJ